OpenStudy (anonymous):

determine the equation of the line wih slope of -3 an containing the point (-7,2)

OpenStudy (angela210793):

y-2=-3(x+7) y=-3x-21+2 y=-3x-19

OpenStudy (aroub):

y=mx+b , m=-3 , (-7,2) 2=-3(-7)+b 2=21+b -21+2=b -19=b so the equation is y=-3x=19